# A Methodological Report on Anomaly Detection on Dynamic Knowledge Graphs

### v2

@@ Abstract
In this paper, we explore different approaches to anomaly detection on dynamic knowledge graphs, specifically in a microservices environment for Kubernetes applications. Our approach explores three dynamic knowledge graph representations: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information. Each phase includes different machine learning and deep learning models. We empirically analyse their performance and propose an approach based on ensemble learning of these models. Our approach significantly outperforms the baseline on the ISWC 2024 Dynamic Knowledge Graph Anomaly Detection dataset, providing a robust solution for anomaly detection in dynamic complex data.
@@ 1 Introduction
Dynamic Knowledge Graphs (DKG), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time. By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
On the leading Micro-Services Architecture (MSA) platform, such as Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms. These data are typically stored in DKGs in the TTL format [^3] https://en.wikipedia.org/wiki/Turtle_(syntax). Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of anomalies.
Despite rapid progress in commercial applications, as a relatively young field, there is still a lack of academic datasets and methodologies. This scarcity can be attributed to several challenges: anomaly events are typically very rare, resulting in an unbalanced data distribution that makes it difficult for models to effectively learn anomaly features and affects their generalisation ability in detecting new anomalies. In addition, models need to capture temporal and structural information and handle heterogeneous nodes/edges, which makes common RNNs and homogeneous GNNs inadequate.
@@
To address these challenges, this paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies. Our method integrates neighbourhood feature extraction with advanced machine learning techniques, including Support Vector Machine (SVM) [6], Extreme Gradient Boosting (XGB) [8], and Self-Attention mechanism (SA) [13] to enhance the model’s ability to capture complex patterns of anomalies. Our approach achieves significant performance improvements over the baseline in the Anomaly Detection on Dynamic Knowledge Graphs challenge (ADDKG) [^4] https://eval.ai/web/challenges/challenge-page/2267/overview at ISWC 2024, providing a robust solution for anomaly detection in dynamic and complex environments.
@@
The structure of the paper is as follows. Section 2 gives the formulation of the problem and describes related work. Section 3 presents the methodology of the paper. Section 4 describes the experimental part of the work and discusses the results obtained. Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
@@ 2.1 Dynamic Knowledge Graphs
In static graphs, common representations include edge lists and adjacency matrices. Among them, the edge list is widely used in software development as it requires less memory. An edge list is a collection of edges, each represented as a pair of nodes.
In the realm of dynamic graphs, existing research categorises them into Continuous-Time Dynamic Graphs (CTDG) and Discrete-Time Dynamic Graphs (DTDG). CTDGs represent dynamic graphs using timestamped events, such as $(u,v,t)$, where $u$ and $v$ are nodes and $t$ is the timestamp of the interaction. In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$ at different time intervals.
Combined with the dynamic graph representations, the Dynamic Knowledge Graph (DKG) is modelled as a heterogeneous dynamic graph, represented by timestamped edge lists. Among the various formats for representing DKGs, Turtle (TTL) is one of the most commonly used. TTL allows the serialisation of nodes, edges, their attributes and timestamps, making it a versatile format for representing entities and relationships in a multi-agent system (MSA). An example of a TTL-based DKG is given in Eq. 1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
$\text{DKG}=\{(s,p,o,t)\mid t\in[t_{1},t_{2},...,T]\}$ (1)
In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$. Table 1 shows all the anomaly types $c$ in the dataset.
@@ 2.2 Anomaly Detection
Traditional anomaly detection methods typically rely on statistical techniques and distance-based measures. However, these methods often struggle with high-dimensional data and complex dependencies [27]. To overcome these problems, common machine learning approaches include Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Isolation Forest (IF) [7]. SVM identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers. XGB is a gradient-boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in separating normal from abnormal data. Isolation Forest is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
@@
With the advancement of software algorithms and hardware, deep learning methods have been increasingly popular for anomaly detection due to their ability to extract complex features [16]. Common architectures for time series include Recurrent Neural Networks (RNN) [9], Temporal Convolutional Networks (TCN) [12], and the Self-Attention mechanism (SA) [13]. RNNs, including Long Short-Term Memory (LSTM) [10] and Gated Recurrent Unit (GRU) [11], capture temporal dependencies, addressing gradient issues in long sequences. TCNs handle long-term dependencies through causal convolutions, while SA captures long-range dependencies by examining relationships across sequences. For graph anomaly detection, Graph Neural Networks (GNN) propagate information through graph structures to identify anomalous patterns [19, 18].
@@
Despite advancements in anomaly detection algorithms, significant challenges persist in dynamic knowledge graphs. Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require large amounts of labelled data, which is often scarce. Although techniques like negative sampling [19, 18, 20] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for dynamic knowledge graphs with attributed edges. Capturing both temporal and structural dependencies remains key to tackling this challenge.
@@ 3 Methodology
This section introduces three representations of dynamic knowledge graphs: sequential data, one-hop, and two-hop graph structures, each adding increasing structural complexity. We then describe the machine learning and deep learning methods applied to these representations, followed by proposed ensemble learning schemes. Figures 1 and 2 provide an overview of the pipeline, including the data structure, the three levels of the data processing, and the final output.
@@
Figure 2: This figure presents the methodology used in this study, with three levels of processing: sequential data, one-hop, and two-hop graph structures. Each level introduces increasing structural complexity, showing how these representations are fed into anomaly detection models.
3.1 DKG as Sequential Data
@@
The time-evolving attributes of nodes can be obtained by extracting raw TTL data, without considering the structural information of the DKG. As shown in the "DKG as Sequential Data" part of Fig. 2, we merge all the sequential data of all the vertices in a DKG by the timestamp. This includes vertices of type $\mathcal{C}$ representing "Cluster", type $\mathcal{P}$ representing "Pod", type $\mathcal{S}$ representing "Service", and type $\mathcal{N}$ representing "Node". For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features. A sliding window then captures temporal dependencies by concatenating the historical information of the past $\tau$ time steps, as shown in Eq. 2.
$\mathbf{f}(t_{j})=[\mathbf{x}(t_{j}-\tau+1)\,|\,\mathbf{x}(t_{j}-\tau+2)\,|\,\cdots\,|\,\mathbf{x}(t_{j})]$ (2)
@@
Therefore, the sequential data representation of a given category $k$ in the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D_{k}}$, where $D_{k}=N_{k}\times d_{k}$, as shown in Eq. 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional information can be computed. For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations.
@@
Based on this data representation, the temporal information across multiple time steps is aggregated into a single vector. Consequently, machine learning algorithms such as SVM, XGB, and IF, as well as deep learning models like MLP, TCN, and RNN, can use this input to capture information across multiple time steps.
3.2 One-Hop Graph Structure
To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities. Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig. 1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp, as shown the aggregate method in the "One-Hop Chart Structure" section of Fig. 2.
$S(t)=\bigcup_{i=1}^{n_{S}}\left(S_{i}(t),\bigcup_{j\in C(S_{i})}\left(C_{j}(t),\bigcup_{k\in P(C_{j})}\left(P_{k}(t),\bigcup_{l\in N(P_{k})}N_{l}(t)\right)\right)\right)$ (4)
For instance, to encode and predict data for the category Service, information related to the Service can be aggregated. Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, Containers and Connections can be aggregated to form the attributes of the Service, as shown in Eq. 4, where $n_{S}$ represents the number of Service instances, $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the elements in the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the elements in the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the elements in the set of Node instances related to $P_{k}$. In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
3.3 Two-Hop Graph Structure
Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model. Inspired by the potential of aggregating second-hop neighbours to provide richer information, we observed that in the tree-like graph of the MSA problem, considering only the current level and its root forms a bipartite graph. For nodes of category $k$, the two-hop neighbours remain nodes of the same category. Recognising potential interdependencies between Service instances at each timestamp, we propose two-hop neighbour aggregation. This can be intuitively achieved by concatenating vertex features, as shown in Eq. 5 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.

### v3

@@ Abstract
In this paper, we explore different approaches to anomaly detection on dynamic knowledge graphs, specifically in a Micro-services environment for Kubernetes applications. Our approach explores three dynamic knowledge graph representations: sequential data, hierarchical data and inter-service dependency data, with each representation incorporating increasingly complex structural information of dynamic knowledge graph. Different machine learning and deep learning models are tested on these representations. We empirically analyse their performance and propose an approach based on ensemble learning of these models. Our approach significantly outperforms the baseline on the ISWC 2024 Dynamic Knowledge Graph Anomaly Detection dataset, providing a robust solution for anomaly detection in dynamic complex data.
@@ 1 Introduction
Dynamic Knowledge Graphs (DKG), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time. By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as Social Media [22], the Internet of Things (IoT) [21, 20], and E-commerce.
Kubernetes-based Micro-services Architecture (MSA) platforms consist of interconnected entities, such as compute nodes, deployable units (pods), and services, which work together to build applications like e-commerce platforms. Since these entities are different and they also depend on each other through different relationships, the best way to represent the data information and the structural information is to construct a Knowledge Graph (KG) [40]. Moreover, the information and dependencies among entities can change over time in MSA platforms. Therefore, the use of a Dynamic Knowledge Graph (DKG) [39] that records a complete snapshot of DKG at each time step is the best way to register and track these changes.
In order to collect and store this DKG dataset, the organisers used the TTL format  [^2] https://en.wikipedia.org/wiki/Turtle_(syntax), a textual syntax for the Resource Description Framework (RDF) that allows KG information to be represented in a compact textual form, such as a triple consisting of a subject, a predicate and an object  [^3] https://www.w3.org/TR/rdf12-turtle/#language-features. Using the Turtle language, the entities and relations of a snapshot of each time step are stored in ternary relations. Predictions based on this structured data is crucial for the operation and optimisation of MSA. This ternary structure representation is similar to the syntax of natural language, which helps to improve the efficiency of querying and parsing, thus improving the effectiveness and robustness of the anomaly detection model.
Although DKGs are well suited for modelling different nodes, variant relationships, and temporal information in Kubernetes-based MSA platforms, their analysing and learning approaches are not well explored yet. In addition, several challenges are specific to the anomaly detection task of Kubernetes-based MSA platforms: 1. Anomalous events are usually very rare, resulting in an imbalanced data distribution, which prevents the model’s ability to learn and generalise anomalous features effectively. 2. Traditional RNNs and homogeneous GNNs are insufficient for accurate anomaly detection due to the need to capture temporal and structural information in Kubernetes environments, coupled with the presence of heterogeneous nodes and edges. To address these challenges, we need to develop specialised approaches capable of handling the dynamic and complex structures of Kubernetes-based DKGs.
@@
This paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies. Our method integrates neighbourhood feature extraction with advanced machine learning techniques, including Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Self-Attention mechanism (SA) [13] to enhance the model’s ability to capture complex patterns of anomalies. Our approach achieves significant performance improvements over the baseline in the Anomaly Detection on Dynamic Knowledge Graphs challenge (ADDKG) [^4] https://eval.ai/web/challenges/challenge-page/2267/overview at ISWC 2024, providing a robust solution for anomaly detection in dynamic and complex environments.
@@
The structure of the paper is as follows. Section 2 gives the problem formulation and describes related work. Section 3 presents the methodology of the paper. Section 4 describes the experimental part of the work and discusses the results obtained. Section 5 concludes our contribution to the ADDKG challenge and discusses future works.
@@ 2.1 Dynamic Knowledge Graphs
In static graphs, common representations include edge lists and adjacency matrices. Among them, the edge list is widely used in software development as it requires less memory. An edge list is a collection of edges. Each edge is represented as a pair of nodes.
In the realm of dynamic graphs, existing research categorises them into Continuous-Time Dynamic Graphs (CTDG) and Discrete-Time Dynamic Graphs (DTDG) [5]. CTDGs represent dynamic graphs using timestamped events, such as $(u,v,t)$, where $u$ and $v$ are nodes and $t$ is the timestamp of the interaction. In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$.
Combined with the dynamic graph representations, the Dynamic Knowledge Graph (DKG) is modelled as a heterogeneous dynamic graph, represented by timestamped edge lists. Among the various formats for representing DKGs, Turtle format (TTL) is used for this competition. The TTL format allows a snapshot of DKGs to be modelled as a series of triples in the structure of (subject, predicate, object). Then a filename that records a timestamp is bound to this snapshot for registering the temporal information. An example of a TTL-based DKG at moment $t$ is given in Eq. 2.1, where $S$ is the ensemble of subject nodes, $O$ is the ensemble of object nodes, $P$ is the ensemble of predicates (or relationships), and $T$ is the ensemble of timestamps.
$\text{DKG}_{t}=\{(s,p,o)\mid s\in S,p\in P,o\in O\}$
Timestamp $\text{Timestamp }t\text{ is associated with DKG}_{t},t\in T$ (1)
In the problem studied in this paper, all anomalies are labelled at the graph level. Specifically, each anomaly in the DKG is labelled with a category $c$ occurring from $t_{start}$ to $t_{end}$.
@@ 2.2 Anomaly Detection
Traditional anomaly detection methods typically rely on statistical techniques and distance-based measures. However, these methods often struggle with high-dimensional data and complex dependencies [26]. To overcome these problems, common machine learning approaches include SVM [6], XGB [8], and Isolation Forest (IF) [7]. SVM identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers. XGB is a gradient-boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in separating normal from abnormal data. IF is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
@@
With the advancement of software algorithms and hardware, deep learning methods have been increasingly popular for anomaly detection due to their ability to extract complex features [15]. Common architectures for time series include Recurrent Neural Networks (RNN) [9], Temporal Convolutional Networks (TCN) [12], and the Self-Attention mechanism (SA) [13]. RNNs, including Long Short-Term Memory (LSTM) [10] and Gated Recurrent Unit (GRU) [11], capture temporal dependencies, addressing gradient issues in long sequences. TCNs handle long-term dependencies through causal convolutions, while SA captures long-range dependencies by examining relationships across sequences.
With the development and popularity of financial, social, traffic and biological networks, graph-based models have gained significant attention in anomaly detection tasks. Graph-based anomaly detection can be classified into four types [27]: node anomalies, edge anomalies, subgraph anomalies and full graph anomalies. Early research [28, 29, 30, 31] focused on static graph anomaly detection, mainly on node anomalies, with limited attention to edge and subgraph anomalies. However, real-world graph data tends to change over time, which introduces additional complexity. To address this problem, researchers have proposed dynamic graph solutions that aim to detect anomalies that occur as the graph changes. Some studies [17, 18, 19] have explored how to detect anomalies at edges or nodes in dynamic graphs. Moreover, networks not only evolve over time, but also present complex structures with heterogeneous nodes and edges. Knowledge Graphs (KGs) provide a powerful framework for modelling such heterogeneous graphs, making them suitable for anomaly detection tasks. Despite the potential of KGs, the application of anomaly detection to (dynamic) KGs is still limited, mainly due to the unstructured nature of KGs and the inherent challenges of dealing with their dynamic and heterogeneous nature.
The growth of cloud services and distributed systems has led to the widespread use of MSA in software and web applications, moving away from traditional monolithic architectures [32]. MSA breaks down an application into independent services, each responsible for a specific function and able to communicate. This architecture enhances the scalability and flexibility of software, but as the number of services increases, it becomes increasingly difficult to manually identify anomalies. Therefore, an automated anomaly detection system is essential to maintain the system.
Most existing work in this area leverages three main data sources (logs [33], performance metrics [34], and distributed traces [35]) and uses machine learning algorithms such as SVMs, K-Nearest Neighbours (KNN) [36], Random Forests (RF) [37], and Multi-Layer Perceptrons (MLPs)[38] for anomaly detection. However, these methods often ignore the structural relationships between services. Modelling MSA as (dynamic) KGs can capture these inter-service dependencies and provide a more comprehensive framework for anomaly detection. However, the integration of DKGs with MSA for anomaly detection is still under-researched, mainly due to the challenges posed by the complexity and scalability of DKGs.
@@
In addition to all these, the anomaly detection task itself presents challenges. Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require a large amount of labelled data and a balance between the amount of normal and abnormal data. However, in practice, the amount of anomalous data tends to be much less than the amount of normal data [27]. Although techniques like negative sampling [18, 17, 19] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for DKGs with attributed edges. Capturing both temporal and structural dependencies remains key to tackling this challenge.
To address these gaps, the ADDKG competition provides a dataset based on MSA and DKGs and encourages participants to develop innovative solutions for anomaly detection in this complex environment.
@@ 3 Methodology
This section introduces three representations of dynamic knowledge graphs: sequential data, hierarchical data, and inter-service dependency data, each adding increasing structural complexity. We then describe the machine learning and deep learning methods applied to these representations, followed by proposed ensemble learning schemes.
3.1 Simplified Tree Structure
@@
The left part of Fig. 1 illustrates the ontology of the DKG utilised in the ADDKG dataset. The ontology comprises a total of ten distinct vertices (also known as functional units in ADDKG DKGs): Cluster, Node, Pod, Container, Image, Connection, Service, Workload, ReplicaSet and Deployment. These vertices are connected to each other with different relationships. Through a comprehensive analysis, three principal observations were identified:
Primary Vertices.
The ontology includes three primary vertices—Node, Container, and Connection. These are designated as primary vertices because only their properties record CPU performance metrics: minimum, maximum, mean, deviation, variance, 50th percentile, 90th percentile, and 95th percentile of CPU resource consumptions at each time step.
Directional Relationships.
All edges within the ontology are unidirectional. For example, relationships such as (Cluster, hasNode, Node) and (Node, hasPod, Pod) demonstrate this unidirectional nature. The only exception is the bidirectional edge pair "controls" and "is controlled by," which essentially represent the same relationship from opposite perspectives.
Simplified Tree Structure.
The DKG can be abstracted into a simplified tree structure, as depicted in the right part of Fig. 1. This simplification consolidates information into three main vertices: Node, Pod, and Service, connected by the relationships (Node, hasPod, Pod) and (Pod, hasConnection, Service). According to the ontology, node attributes from the primary vertices (Container and Connection) can be effectively propagated to Pod and Service through the "hasContainer" and "hasConnection" relations.
The illustrated tree structure in Fig. 1, comprising Node, Pod, and Service, serves as a reduced version of the tree structure, hence it can not represent the whole DKG. We will use this example to facilitate the later explanation of the three data representation methods used in our study. In the example, yellow vertices represent Nodes, pink vertices represent Pods, and green vertices represent Services. Each vertex type can have varying numbers of instances; for instance, this example includes one Node (n1), three Pods (p1, p2, p3), and four Services (s1, s2, s3, s4).
3.2 Sequential Data Representation
In this phase, the representation of the DKGs can be constructed by concatenating the time-evolving attributes of the vertices in the simplified tree structure, specifically from vertices Node, Pod and Service.
Figure 2: This figure presents the sequential data representation based on the simplified tree structure.
Operationally, time-evolving attributes can be extracted by querying the nodes and relations from the triples in TTL files. We can then simply merge these data by timestamps to form the sequential data representation. This operation largely ignores the structural relations between nodes, e.g., (Pod, hasNode, Node).
As shown in the "Sequential Data Representation" part of Fig. 2, temporal data from Node 1 is systematically merged into Pods 1, 2, and 3 based on their respective timestamps. Then the merged data from these Pods is further consolidated into Services 1, 2, 3, and 4, ensuring temporal consistency throughout the hierarchical aggregation process.
@@ 3.2 Sequential Data Representation
Mathematically, for each given vertex category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{T\times D_{k}}$, where $T$ indicates the number of snapshots, and $D_{k}$ indicates the number of features of category $k$ by concatenating all the instances, which equals to $N_{k}\times d_{k}$ with $N_{k}$ indicates the number of instances of the category $k$, and $d_{k}$ is the number of features of the category $k$. Then we can merge all vertices Node, Pod and Service together based on timestamps, thus we can get a $\mathbf{\tilde{X}}\in\mathbb{R}^{T\times D}$, where $D=D_{Node}+D_{Pod}+D_{Service}=\sum^{3}_{k=1}N_{k}*d_{k}$. A sliding window then captures temporal dependencies by concatenating the historical information of the past H time steps, as shown in Eq. 2.
$\mathbf{f}(t_{j})=[\mathbf{\tilde{x}}(t_{j}-H+1)\,|\,\mathbf{\tilde{x}}(t_{j}-H+2)\,|\,\cdots\,|\,\mathbf{\tilde{x}}(t_{j})]$ (2)
@@
Therefore, the sequential data representation the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D\times H}$, where $H$ is the number of historical steps, as shown in Eq. 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{\tilde{x}}(t_{1}-H+1)&\mathbf{\tilde{x}}(t_{1}-H+2)&\cdots&\mathbf{\tilde{x}}(t_{1})\\ \mathbf{\tilde{x}}(t_{2}-H+1)&\mathbf{\tilde{x}}(t_{2}-H+2)&\cdots&\mathbf{\tilde{x}}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{\tilde{x}}(t_{m}-H+1)&\mathbf{\tilde{x}}(t_{m}-H+2)&\cdots&\mathbf{\tilde{x}}(t_{m})\end{bmatrix}$ (3)
Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional time-dependent features can be computed to enhance the model’s ability to detect anomalies. For example, calculating the difference between consecutive data points can capture the trend of the time series and understand the direction of change. Similarly calculating the variance of data within a specific time window can reflect the stability of the data. By additionally incorporating the difference and variance as temporal features, the dynamic behaviour of the system can be effectively captured and the ability to identify anomalies can be improved.
@@
Based on this data representation, machine learning algorithms such as SVM, XGB, and IF, as well as deep learning models like MLP, TCN, LSTM, GRU and SA, can use this input to capture information across multiple time steps.
3.3 Hierarchical Data Representation
Figure 3: This figure presents the hierarchical data representation based on the simplified tree structure.
To investigate the structural information within the DKGs, we propose a hierarchical neighbourhood information aggregation method designed to capture the connectivity relationships among various entities. In the simplified tree structure, hierarchical relationships exist between vertices. For example, a Node has multiple Pods and a Pod is connected to different Services through multiple Connections and the "hasConnection" relationship (see Fig. 1).
Operationally, when relationships exist across Nodes, Pods, and Services, we can progressively aggregate their instances’ time-evolving attributes from Nodes to Services using aggregation methods such as mean, max, and min. This approach enables the construction of a structural-temporal representation for each Service instance at each timestamp, as illustrated in the "Hierarchical Data Structure" in Fig. 3.
This operation influences the feature dimension $D$ of $\mathbf{X}$, defined in Section 3.2. Initially, $D$ was determined by $N_{k}$, the number of instances of category $k$. However, we have now redefined $D$ into $\sum^{3}_{k=1}AGG*d_{k}$, with $AGG$ corresponding to the number of aggregation methods employed.
This aggregation process is analogous to a multi-hop operation, where information is aggregated from multiple hops away from the target Service. For instance, for Service 1 Fig. 3, temporal data is first aggregated from Node 1, followed by the aggregation of data from Pods 1, 2, and 3, and finally merging them based on the timestamp. This method ensures that the temporal dynamics and structural relationships inherent in the DKG are effectively captured, thereby enhancing the accuracy and robustness of anomaly detection within the MSA.
3.4 Hierarchical and Inter-Service Representation
Some graph machine learning algorithms have pointed out that a larger convolutional field can optimise the model [5]. Since we have already aggregated the multi-hop structural-temporal information into the Service vertex, in this section we supposed that all Service instances can influence each other. Therefore, we can further enrich the DKG representations by providing additional correlations between Service instances.
Figure 4: This figure presents the inter-service data representation based on the simplified tree structure.
This representation can be intuitively achieved by concatenating vertex features, as shown in Eq. 4 where $N_{\mathcal{S}}$ represents the number of Service instances and $\mathbf{s}_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$. However, this approach will greatly increase the size of the feature dimension $D$ of $\mathbf{X}$, requiring more computational resources during feature processing and model training.